# Animatable Gaussians: Learning Pose-dependent Gaussian Maps for High-fidelity Human Avatar Modeling

### v3

@@ Abstract
Modeling animatable human avatars from RGB videos is a long-standing and challenging problem. Recent works usually adopt MLP-based neural radiance fields (NeRF) to represent 3D humans, but it remains difficult for pure MLPs to regress pose-dependent garment details. To this end, we introduce Animatable Gaussians, a new avatar representation that leverages powerful 2D CNNs and 3D Gaussian splatting to create high-fidelity avatars. To associate 3D Gaussians with the animatable avatar, we learn a parametric template from the input videos, and then parameterize the template on two front & back canonical Gaussian maps where each pixel represents a 3D Gaussian. The learned template is adaptive to the wearing garments for modeling looser clothes like dresses. Such template-guided 2D parameterization enables us to employ a powerful StyleGAN-based CNN to learn the pose-dependent Gaussian maps for modeling detailed dynamic appearances. Furthermore, we introduce a pose projection strategy for better generalization given novel poses. Overall, our method can create lifelike avatars with dynamic, realistic and generalized appearances. Experiments show that our method outperforms other state-of-the-art approaches.
Figure 1: Lifelike animatable avatars with highly dynamic, realistic and generalized details created by our method.
footnotetext: Corresponding author: Yebin Liu.
1 Introduction
@@
Overall, benefiting from these contributions, our method can create lifelike animatable avatars with highly dynamic, realistic and generalized appearances as shown in Fig. 1.
2 Related Work
2.1 Mesh-based Human Avatars
@@
Figure 2: Illustration of the pipeline. It contains two main steps: 1) Reconstruct a character-specific template from multi-view images. 2) Predict pose-dependent Gaussian maps through the StyleUNet, and render the synthesized avatar by LBS and differentiable rasterization.
2.2 Implicit Function-based Human Avatars
@@
2.3 Point-based Human Avatars
@@
Point cloud is also a powerful and popular representation in human avatar modeling. Given 3D scans of a character, SCALE [51] and POP [52] learn the non-rigid deformation of dense points on SMPL UV maps to represent the dynamic garment wrinkles. FITE [46] and CloSET [100] extract pose features from projective maps or PointNet [65, 66] to avoid discontinuity on the UV map. SKiRT [53] and FITE learn a coarse template from the input scans and utilize learned or diffused skinning weights to animate loose clothes. Prokudin et al. [64] propose dynamic point fields for general dynamic reconstruction. This work and NPC [78] show results on avatars created from RGB videos using Point-NeRF [93]. However, applying Point-NeRF to avatar modeling still relies on a low-frequency coordinate-based MLP, struggling with the same problems in Sec. 2.2. On the other hand, point-based rendering via splatting [63, 109, 110, 111, 99, 1, 41, 39, 71] offers another probability for animatable avatar modeling. PointAvatar [105] learns a canonical point cloud and deformation field to model head avatars from a monocular video via PyTorch3D’s [69] differentiable point renderer. Recently, 3D Gaussian splatting (3DGS) [35], an efficient differentiable point-based rendering method, has been proposed for real-time photo-realistic scene rendering. In addition to applying 3DGS to dynamic scenes [50, 96, 89, 97], human reconstruction [104], editing [10, 73] and talking heads [67, 94], many concurrent works also introduce 3DGS into animatable human avatars. However, most of them [108, 42, 68, 27, 38] adopt MLPs to regress the Gaussian attributes, leading to blurry avatar appearances. ASH [58] and GaussianAvatar [26] parameterize the 3D character on a 2D UV map to predict Gaussian attributes using 2D CNNs, sharing similar ideas with us.
3 Method
3.1 Preliminary: 3D Gaussian Splatting
@@ 3.2 Overview
    Learning Pose-dependent Gaussian Maps. Given a training pose, we first deform the template to the posed space via linear blend skinning (LBS) and render the posed vertex coordinates to canonical front & back views to obtain two position maps. The position maps serve as the pose condition and are translated into front & back Gaussian maps through a StyleUNet [83]. We then extract valid 3D Gaussians inside the template mask, and deform the canonical 3D Gaussians to the posed space by LBS. Eventually, we render the posed 3D Gaussians to a given camera view through differentiable splatting-based rasterization [35].
Figure 3: Illustration of the posed position maps.
3.3 Avatar Representation
Learning Parametric Template. Given the multi-view videos, we first select one frame in which the character is under a near A-pose. Our goal is to reconstruct a canonical geometric model as the template from the multi-view images. Specifically, we represent the canonical character as an SDF and color field instantiated by an MLP. To associate the canonical and posed spaces, we precompute a skinning weight volume $\mathcal{W}$ in the canonical space by diffusing the weights from the SMPL surface throughout the whole 3D volume along the surface normal [46]. For each point in the posed space, we search its canonical correspondence by root finding [7]:
@@ 3.3 Avatar Representation
Template-guided Parameterization. Previous human avatar representations in NeRF-based approaches [61, 47, 106] necessitate the coordinate-based MLPs for the formulation of the implicit NeRF function. However, MLPs have demonstrated a low-frequency bias [80], hindering their ability to model high-frequency human dynamics. In light of this observation, we replace MLPs with more powerful 2D CNNs for creating higher-quality human avatars. To ensure compatibility with 2D networks, the 3D representation of the human avatar needs to be parameterized in 2D space. Therefore, we propose to parameterize the 3D Gaussians anchored on the canonical template onto front & back views via orthogonal projection. As illustrated in Fig. 3, given a driving pose $\mathbf{\Theta}$, we first deform the template to the posed space via LBS. Note that we do not consider the global transformation in this skinning process, because the global orientation and translation would not change the human dynamic details. Then we take the posed coordinate as the vertex color on the canonical template, and render it to both front & back views by orthogonal projection, obtaining posed position maps $\mathcal{P}_{\text{f}}(\mathbf{\Theta})$ and $\mathcal{P}_{\text{b}}(\mathbf{\Theta})$ that serve as pose conditions for the network.
@@
Pose-dependent Gaussian Maps. We employ a powerful StyleGAN-based CNN , StyleUNet [83] $\mathcal{F}$, to predict pose-dependent Gaussian maps from the pose conditions:
$\mathcal{G}_{\text{f}}(\mathbf{\Theta}),\mathcal{G}_{\text{b}}(\mathbf{\Theta})\leftarrow\mathcal{F}(\mathcal{P}_{\text{f}}(\mathbf{\Theta}),\mathcal{P}_{\text{b}}(\mathbf{\Theta}),\mathcal{V}),$ (4)
@@
where $\mathcal{G}_{\text{f}}(\mathbf{\Theta})$ and $\mathcal{G}_{\text{b}}(\mathbf{\Theta})$ are front and back pose-dependent Gaussian maps, respectively, and each pixel represents a 3D Gaussian [35] including a position, covariance, opacity and color. We also modulate the output color attributes on Gaussian maps with a view direction map $\mathcal{V}$ to model view-dependent variance like NeRF-based approaches [61]. We extract canonical 3D Gaussians inside the template mask from the pose-dependent Gaussian maps. It is worth mentioning that despite utilizing only front and back views for parameterizing the 3D Gaussians, the resulting point clouds still cover the side regions and hands of the human body as demonstrated in Fig. 4. The reason is that the projection to front & back views is orthographic, thus there exist sufficient 3D Gaussians to model these parts.
Figure 4: Canonical 3D Gaussians on side regions and hands.
@@
LBS of 3D Gaussians. To render the synthesized avatar under the driving pose, we need to deform the canonical 3D Gaussians to the posed space. Specifically, given a canonical 3D Gaussian, we transform its position $\mathbf{p}_{\text{c}}$ and covariance $\mathbf{\Sigma}_{\text{c}}$ attributes:
@@
Training. To ensure that the position attribute of predicted Gaussian maps approximates the canonical human body, we opt to predict an offset map $\Delta\mathcal{O}(\mathbf{\Theta})$ on the parametric template instead of a position map. Our training losses include L1 and perceptual losses [101] between the rendered images and ground truth, and a regularization loss:
$\mathcal{L}=\mathcal{L}_{1}+\lambda_{\text{perceptual}}\mathcal{L}_{\text{perceptual}}+\lambda_{\text{reg}}\mathcal{L}_{\text{reg}},$ (6)
where $\lambda$s are loss weights, and the regularization loss $\mathcal{L}_{\text{reg}}=\|\Delta\mathcal{O}(\mathbf{\Theta})\|_{2}^{2}$ restrains the predicted offsets from being extremely large.
3.4 Pose Projection Strategy

### v4

# Animatable and Relightable Gaussians for High-fidelity Human Avatar Modeling

@@ Abstract
Modeling animatable human avatars from RGB videos is a long-standing and challenging problem. Recent works usually adopt MLP-based neural radiance fields (NeRF) to represent 3D humans, but it remains difficult for pure MLPs to regress pose-dependent garment details. To this end, we introduce Animatable Gaussians, a new avatar representation that leverages powerful 2D CNNs and 3D Gaussian splatting to create high-fidelity avatars. To associate 3D Gaussians with the animatable avatar, we learn a parametric template from the input videos, and then parameterize the template on two front & back canonical Gaussian maps where each pixel represents a 3D Gaussian. The learned template is adaptive to the wearing garments for modeling looser clothes like dresses. Such template-guided 2D parameterization enables us to employ a powerful StyleGAN-based CNN to learn the pose-dependent Gaussian maps for modeling detailed dynamic appearances. Furthermore, we introduce a pose projection strategy for better generalization given novel poses. To tackle the realistic relighting of animatable avatars, we introduce physically-based rendering into the avatar representation for decomposing avatar materials and environment illumination. Overall, our method can create lifelike avatars with dynamic, realistic, generalized and relightable appearances. Experiments show that our method outperforms other state-of-the-art approaches.
Index Terms:
Animatable avatar, human reconstruction, view synthesis, animation, relighting.
Figure 1: Lifelike relightable and animatable avatars with highly dynamic, realistic and generalized details created by our method. We show synthesized results animated by the same pose under the capture environment and novel lights.
I Introduction
@@
A preliminary version of this work has been published in CVPR 2024 [18], in which we propose a novel avatar representation for modeling realistic animatable human avatars from multi-view videos. However, the preliminary work [18] can only animate the avatar under illumination from the capture environment. In the current version, we further introduce physically-based rendering (PBR) [19, 20] into our avatar representation for creating both animatable and relightable human avatars. Specifically, besides the original attributes of 3D Gaussians, our model also predicts albedo, roughness, and light visibility to decompose the avatar materials and illumination of the capture environment. As a result, our model can produce realistic animation under novel illumination.
We extend the preliminary version [18] in the following ways. First, we introduce PBR into our avatar representation (Animatable Gaussians) for creating relightable human avatars. Second, we compare our method with concurrent works on 3D Gaussian splatting-based avatars, and our method outperforms them on the avatar quality. Third, we compare our method with other works on human performance relighting, and our method can produce more realistic human relighting.
@@
  • We introduce physically-based rendering into Animatable Gaussians for photorealistic relighting under novel illumination.
Overall, benefiting from these contributions, our method can create lifelike animatable and relightable avatars with highly dynamic, realistic and generalized appearances as shown in Fig. 1. The code is available at https://github.com/lizhe00/AnimatableGaussians, and earns more than 700 stars.
II Related Work
II-A Mesh-based Human Avatars
@@
II-B Implicit Function-based Human Avatars
@@
II-C Point-based Human Avatars
@@
Point cloud is also a powerful and popular representation in human avatar modeling. Given 3D scans of a character, SCALE [67] and POP [3] learn the non-rigid deformation of dense points on SMPL UV maps to represent the dynamic garment wrinkles. FITE [11] and CloSET [68] extract pose features from projective maps or PointNet [69, 70] to avoid discontinuity on the UV map. SKiRT [12] and FITE learn a coarse template from the input scans and utilize learned or diffused skinning weights to animate loose clothes. Prokudin et al.[71] propose dynamic point fields for general dynamic reconstruction. This work and NPC [72] show results on avatars created from RGB videos using Point-NeRF [73]. However, applying Point-NeRF to avatar modeling still relies on a low-frequency coordinate-based MLP, struggling with the same problems in Sec. II-B. On the other hand, point-based rendering via splatting [74, 75, 76, 77, 78, 79, 80, 81, 82] offers another probability for animatable avatar modeling. PointAvatar [83] learns a canonical point cloud and deformation field to model head avatars from a monocular video via PyTorch3D’s [84] differentiable point renderer.
Recently, 3D Gaussian splatting [10] (3DGS), an efficient differentiable point-based rendering method, has been proposed for real-time photo-realistic scene rendering. Along with extending 3DGS to dynamic scene modeling [85, 86, 87, 88], many researchers also introduced 3DGS into animatable human avatars. Specifically, D3GA [89] leverages cage-based deformation to model the motion of 3D Gaussians. While other approaches like GART[90], 3DGS-Avatar [91], GauHuman [92] and HUGS [93] employ linear blend skinning (LBS) to model human motions and reconstruct a 3DGS-based animatable avatar from monocular videos. However, these approaches cannot produce highly realistic and dynamic human appearances because they all represent the canonical 3D human using MLPs, facing the same low-frequency problem as NeRF-based approaches (Sec. II-B). We observe such an explicit point-based representation can be combined with 2D CNNs for high-quality avatar modeling. Concurrent works including ASH [94] and GaussianAvatar [95] parameterize the 3D character on a 2D UV map to predict Gaussian attributes using 2D CNNs, sharing similar ideas with our method. Differently, we parameterize the canonical 3D human on the front and back views by orthographic projection.
II-D Human Relighting
Human relighting aims to manipulate the reflectance field of the human surface, thus enabling an immersive fusion with novel illumination. Conventional approaches [96, 97, 98, 99, 100, 101, 102, 103] propose capturing the reflectance characteristics of a human subject through a LightStage arrangement. This setup entails controlled illumination systems and dense camera arrays, facilitating the generation of photorealistic renderings under diverse lighting conditions. However, such configurations are both financially demanding to capture and not readily accessible to the public. With the advancement of neural implicit representations, recent methods [104, 105, 106, 36, 107, 108] necessitate solely multi-view or even monocular video recordings obtained under constant unknown illumination conditions to model both human motion and light transport properties. Relighting4D [104] employs NeuralBody [109] as the dynamic human model and utilizes neural inverse rendering to decompose it into a 3D reflectance field, enabling the estimation of materials and lighting properties. Building upon this framework, Sun et al.[106] apply LBS to transform the reflectance field from the canonical space to the observation space, facilitating animated sequences with relighting effects. However, this work does not address shadowing effects. RANA [105] pretrains an SMPL+D-based representation incorporating albedo and normal map refinement techniques. It utilizes a simplified spherical harmonics lighting model to achieve relighting effects but lacks the capability to accurately model specular effects and shadows. Lin et al.[107] propose to estimate pose-aware light visibility through part-wise MLPs, demonstrating improved shadowing estimation when generalizing to unseen poses. Xu et al.[36] utilize Hierarchical Distance Queries (HDQ) via sphere tracing to calculate correct SDF values under arbitrary human poses and then incorporate distant field soft shadow (DFSS) for estimating reasonable soft visibility maps. IntrinsicAvatar [108] employs explicit Monte-Carlo ray tracing in canonical space to capture secondary shading effects, thereby enabling precise estimation of materials and environmental lighting. However, most previous methods rely on NeRF-based techniques, which inherently yield limited rendering quality. Thanks to the proposed powerful Gaussian representation, our method can achieve more vivid and realistic relighting results under novel poses.
Figure 2: Illustration of the avatar modeling pipeline. It contains two main steps: 1) Reconstruct a character-specific template from multi-view images. 2) Predict pose-dependent Gaussian and intrinsic maps through StyleUNets, and render the posed Gaussians by Gaussian splatting and physically-based rendering to learn both pose-dependent dynamics and avatar materials. Finally, given a novel environment light, we can animate the avatar with realistic dynamic appearances and shadow effects.
III Method
III-A Preliminary: 3D Gaussian Splatting
@@ III-B Overview
    Learning Pose-dependent Gaussian and Intrinsic Maps. Given a training pose, we first deform the template to the posed space via linear blend skinning (LBS) and render the posed vertex coordinates to canonical front & back views to obtain two position maps. The position maps serve as the pose condition and are translated into front & back Gaussian and intrinsic maps through StyleUNets [17]. We then extract valid 3D Gaussians inside the template mask from the Gaussian map, and deform the canonical 3D Gaussians to the posed space by LBS. In one branch, we directly render the posed 3D Gaussians to a camera view using the observed colors on the Gaussian maps. On the other hand, we render the posed 3D Gaussians using the PBR color computed from the intrinsic maps. These two branches allow our method not only to learn pose-dependent animation but also to decompose the avatar appearance into materials and light conditions for relighting purposes.
III-C Avatar Representation
III-C1 Learning Parametric Template
Given the multi-view videos, we first select one frame in which the character is under a near A-pose. Our goal is to reconstruct a canonical geometric model as the template from the multi-view images. Specifically, we represent the canonical character as an SDF and color field instantiated by an MLP. To associate the canonical and posed spaces, we precompute a skinning weight volume $\mathcal{W}$ in the canonical space by diffusing the weights from the SMPL surface throughout the whole 3D volume along the surface normal [11]. For each point in the posed space, we search its canonical correspondence by root finding [44]:
@@ III-C1 Learning Parametric Template
Figure 3: Illustration of the posed position maps.
III-C2 Template-guided Parameterization
@@
Previous human avatar representations in NeRF-based approaches [5, 6, 7] necessitate the coordinate-based MLPs for the formulation of the implicit NeRF function. However, MLPs have demonstrated a low-frequency bias [9], hindering their ability to model high-frequency human dynamics. In light of this observation, we replace MLPs with more powerful 2D CNNs for creating higher-quality human avatars. To ensure compatibility with 2D networks, the 3D representation of the human avatar needs to be parameterized in 2D space. Therefore, we propose to parameterize the 3D Gaussians anchored on the canonical template onto front & back views via orthogonal projection. As illustrated in Fig. 3, given a driving pose $\mathbf{\Theta}$, we first deform the template to the posed space via LBS. Note that we do not consider the global transformation in this skinning process, because the global orientation and translation would not change the human dynamic details. Then we take the posed coordinate as the vertex color on the canonical template, and render it to both front & back views by orthogonal projection, obtaining posed position maps $\mathcal{P}_{\text{f}}(\mathbf{\Theta})$ and $\mathcal{P}_{\text{b}}(\mathbf{\Theta})$ that serve as pose conditions for the network.
III-C3 Pose-dependent Gaussian Maps
@@
We employ a powerful StyleGAN-based CNN , StyleUNet [17] $\mathcal{F}_{\mathcal{G}}$, to predict pose-dependent Gaussian maps from the pose conditions:
$\mathcal{G}_{\text{f}}(\mathbf{\Theta}),\mathcal{G}_{\text{b}}(\mathbf{\Theta})\leftarrow\mathcal{F}_{\mathcal{G}}(\mathcal{P}_{\text{f}}(\mathbf{\Theta}),\mathcal{P}_{\text{b}}(\mathbf{\Theta}),\mathcal{V}),$ (4)
@@
where $\mathcal{G}_{\text{f}}(\mathbf{\Theta})$ and $\mathcal{G}_{\text{b}}(\mathbf{\Theta})$ are front and back pose-dependent Gaussian maps, respectively, and each pixel represents a 3D Gaussian [10] including a position, covariance, opacity and color. To ensure that the position attribute of predicted Gaussian maps approximates the canonical human body, we opt to predict an offset map $\Delta\mathcal{O}(\mathbf{\Theta})$ on the parametric template instead of a global position map. We also modulate the output color attributes on Gaussian maps with a view direction map $\mathcal{V}$ to model view-dependent variance like NeRF-based approaches [5]. We extract canonical 3D Gaussians inside the template mask from the pose-dependent Gaussian maps. It is worth mentioning that despite utilizing only front and back views for parameterizing the 3D Gaussians, the resulting point clouds still cover the side regions and hands of the human body as demonstrated in Fig. 4. The reason is that the projection to front & back views is orthographic, thus there exist sufficient 3D Gaussians to model these parts.
III-C4 LBS of 3D Gaussians
@@
To render the synthesized avatar under the driving pose, we need to deform the canonical 3D Gaussians to the posed space. Specifically, given a canonical 3D Gaussian, we transform its position $\mathbf{p}_{\text{c}}$ and covariance $\mathbf{\Sigma}_{\text{c}}$ attributes:
@@
Figure 4: Canonical 3D Gaussians on side regions and hands.
III-C5 Pose-dependent Intrinsic Maps
The pose-dependent Gaussian maps only represent the human appearances under the illumination of the capture environment, limiting animation under novel lighting conditions. To disentangle the avatar geometry, material and lighting conditions, we leverage the classic rendering equation [112] to simulate the rendering process:
[…]
where $L_{o}(\mathbf{x},\boldsymbol{\omega}_{o})$ and $L_{i}(\mathbf{x},\boldsymbol{\omega}_{i})$ are the outgoing and incident radiance at a surface position $\mathbf{x}$ along direction $\boldsymbol{\omega}_{o}$ and $\boldsymbol{\omega}_{i}$, respectively, and $\mathbf{n}$ is the normal vector at $\mathbf{x}$. $f(\mathbf{x},\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o};\alpha,\gamma,\mathbf{n})$ is the Bidirectional Reflectance Distribution Function (BRDF) determined by the surface geometry (normal $\mathbf{n}$) and material properties including albedo $\alpha$ and roughness $\gamma$. Considering the light visibility, the incident radiance is further formulated as
$L_{\text{i}}(\mathbf{x},\boldsymbol{\omega}_{i})=V(\mathbf{x},\boldsymbol{\omega}_{i})L_{i}(\boldsymbol{\omega}_{i}),$ (7)
where $V(\mathbf{x},\boldsymbol{\omega}_{i})\in\{0,1\}$ indicates whether $\mathbf{x}$ is visible along the light direction $\boldsymbol{\omega}_{i}$. The global light $L_{i}(\boldsymbol{\omega}_{i})$ is parametrized as learnable Spherical Harmonics (SH), which are optimized during the inverse rendering process.
Based on the physically-based rendering process, we additionally learn pose-dependent intrinsic maps including normal $\mathcal{N}(\mathbf{\Theta})$, albedo $\mathcal{A}(\mathbf{\Theta})$, and roughness $\boldsymbol{\gamma}(\mathbf{\Theta})$ maps on the front and back canonical views:
$\begin{split}\mathcal{N}_{\text{f}}(\mathbf{\Theta}),\mathcal{N}_{\text{b}}(\mathbf{\Theta})\leftarrow\mathcal{F}_{\mathcal{N}}(\mathcal{P}_{\text{f}}(\mathbf{\Theta}),\mathcal{P}_{\text{b}}(\mathbf{\Theta})),\\ \mathcal{A}_{\text{f}}(\mathbf{\Theta}),\mathcal{A}_{\text{b}}(\mathbf{\Theta})\leftarrow\mathcal{F}_{\mathcal{A}}(\mathcal{P}_{\text{f}}(\mathbf{\Theta}),\mathcal{P}_{\text{b}}(\mathbf{\Theta})),\\ \boldsymbol{\gamma}_{\text{f}}(\mathbf{\Theta}),\boldsymbol{\gamma}_{\text{b}}(\mathbf{\Theta})\leftarrow\mathcal{F}_{\boldsymbol{\gamma}}(\mathcal{P}_{\text{f}}(\mathbf{\Theta}),\mathcal{P}_{\text{b}}(\mathbf{\Theta})),\end{split}$ (8)
where $\mathcal{F}_{\mathcal{N}}$, $\mathcal{F}_{\mathcal{A}}$ and $\mathcal{F}_{\boldsymbol{\gamma}}$ are StyleUNet modules. More specifically, the learned normal map $\mathcal{N}(\mathbf{\Theta})$ is an offset of the surface normal on the parametric template for better pose generalization. The light visibility $V(\mathbf{x},\boldsymbol{\omega}_{i})$ can be computed from the posed 3D Gaussians by point-based ray tracing [20]. However, as mentioned in [20], computing the light visibility during training is not preferred because of the computational complexity. Therefore, we train an additional network to predict light visibility and supervise the prediction by randomly sampling view directions. Specifically, we formulate the light-direction-dependent visibility as SH, and predict SH coefficient map $\mathcal{K}(\mathbf{\Theta})$ using a StyleUNet $\mathcal{F}_{\mathcal{K}}$:
$\mathcal{K}_{\text{f}}(\mathbf{\Theta}),\mathcal{K}_{\text{b}}(\mathbf{\Theta})\leftarrow\mathcal{F}_{\mathcal{K}}(\mathcal{P}_{\text{f}}(\mathbf{\Theta}),\mathcal{P}_{\text{b}}(\mathbf{\Theta})).$ (9)
Similar to the Gaussian maps, each pixel value on normal, albedo, roughness and light visibility maps is associated with a 3D Gaussian. Following Relightable 3D Gaussian [20], we sample $N$ incident light directions over the hemisphere space, and calculate the PBR color of each 3D Gaussian by the discrete form of Eq. 6:
$\begin{split}&L_{o}(\boldsymbol{\mu},\boldsymbol{\omega}_{o})=\\ &\sum_{\boldsymbol{\omega}_{i}}L_{i}(\boldsymbol{\mu},\boldsymbol{\omega})f(\boldsymbol{\mu},\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o};\alpha_{\boldsymbol{\mu}},\gamma_{\boldsymbol{\mu}},\mathbf{n}_{\boldsymbol{\mu}})(\boldsymbol{\omega}_{i}\cdot\mathbf{n}_{\mu})\Delta\boldsymbol{\omega}_{i},\end{split}$ (10)
where ${\boldsymbol{\mu}}$ is the position of a 3D Gaussian, $\alpha_{\boldsymbol{\mu}}$, $\gamma_{\boldsymbol{\mu}}$ and $\mathbf{n}_{\boldsymbol{\mu}}$ is the corresponding albedo, roughness and normal attribute. Finally, the physically-based rendered image can be obtained by rasterization (Eq. 2).
III-D Training
The optimizable parameters of the avatar model include the parameters of StyleUNets and an environment light map $\boldsymbol{l}_{\text{env}}$. Our training loss consists of five parts: the reconstruction loss $\mathcal{L}_{\text{recon}}$, PBR loss $\mathcal{L}_{\text{PBR}}$, normal loss $\mathcal{L}_{\text{normal}}$, visibility loss $\mathcal{L}_{\text{vis}}$ and regularization loss $\mathcal{L}_{\text{reg}}$:
$\mathcal{L}=\mathcal{L}_{\text{recon}}+\mathcal{L}_{\text{PBR}}+\lambda_{\text{normal}}\mathcal{L}_{\text{normal}}+\lambda_{\text{vis}}\mathcal{L}_{\text{vis}}+\mathcal{L}_{\text{reg}}.$ (11)
III-D1 Reconstruction Loss
The reconstruction loss aims to learn the geometry and the observed texture of the human avatar from the multi-view videos without considering the PBR process. We denote the rendered image using color attributes from the Gaussian maps as $\mathbf{C}_{\text{OBS}}$. The reconstruction loss involves an L1 loss and a perceptual loss [113] between $\mathbf{C}_{\text{OBS}}$ and the ground-truth image $\mathbf{C}_{\text{GT}}$:
$\mathcal{L}_{\text{recon}}=\mathcal{L}_{1}(\mathbf{C}_{\text{OBS}},\mathbf{C}_{\text{GT}})+\lambda_{\text{perceptual}}\mathcal{L}_{\text{perceptual}}(\mathbf{C}_{\text{OBS}},\mathbf{C}_{\text{GT}}).$ (12)
III-D2 PBR Loss
The PBR loss aims to disentangle the light condition and avatar materials from the multi-view observations. We denote the rendered image by the PBR process (i.e., Eq. 10) as $\mathbf{C}_{\text{PBR}}$. The PBR loss involves an L1 loss and a perceptual loss between $\mathbf{C}_{\text{PBR}}$ and $\mathbf{C}_{\text{GT}}$:
$\mathcal{L}_{\text{PBR}}=\mathcal{L}_{1}(\mathbf{C}_{\text{PBR}},\mathbf{C}_{\text{GT}})+\lambda_{\text{perceptual}}\mathcal{L}_{\text{perceptual}}(\mathbf{C}_{\text{PBR}},\mathbf{C}_{\text{GT}}).$ (13)
III-D3 Normal Loss
We employ a pretrained normal estimation network [114] to supervise our predicted normal. Specifically, we render a normal image with the predicted normal as additional channels by rasterization, and compare it with the estimated one by [114] using L1 loss.
III-D4 Visibility Loss
In each iteration, for each 3D Gaussian, we randomly sample a light direction over the hemisphere around its normal. We compute L1 loss between the predicted visibility of the sampled directions and the ground-truth one obtained by point-based ray tracing[20].
III-D5 Regularization Loss
For the stability and convergence of avatar training, we design several regularization losses on the geometry and materials. Specifically, the regularization losses involve a geometric regularization loss and smooth regularization losses on albedo and roughness:
$\mathcal{L}_{\text{reg}}=\lambda_{\text{geo}}\mathcal{L}_{\text{geo}}+\lambda_{\text{albedo}}\mathcal{L}_{\text{smooth}}(\mathbf{I}_{\mathcal{A}})+\lambda_{\text{roughness}}\mathcal{L}_{\text{smooth}}(\mathbf{I}_{\boldsymbol{\gamma}}),$ (14)
where $\mathcal{L}_{\text{geo}}=\|\Delta\mathcal{O}\|_{2}^{2}$ restrains the predicted offset map $\Delta\mathcal{O}$ from being extremely large, and the smooth loss is a bilateral smoothness term [20] that constrains the material properties to change continuously in areas with smooth colors:
$\mathcal{L}_{\text{smooth}}(\mathbf{I})=\|\nabla\mathbf{I}\|\exp(-\|\nabla\mathbf{C}_{\text{GT}}\|),$ (15)
where $\mathbf{I}$ is the rendered albedo or roughness map ($\mathbf{I}_{\mathcal{A}}$ or $\mathbf{I}_{\boldsymbol{\gamma}}$) via splatting-based rasterization.
III-E Animation and Relighting
III-E1 Pose Projection Strategy